OpenStudy (anonymous):

HELP I WILL GIVE A MEDAL AND I WILL FAN!!! A system of equations is shown below: x + y = 3 2x – y = 6 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (solomonzelman):

add the equations to each other.

OpenStudy (anonymous):

so like how do i do that like im confused

OpenStudy (solomonzelman):

x + y = 3 2x – y = 6 ---------------- 3x + 0y = 9 good ?

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

so like what else do I do @SolomonZelman

OpenStudy (solomonzelman):

3x+0y=9 is the result after the addition./ Can you tell me what x equals ?

OpenStudy (anonymous):

how do I find what it equals

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

actually wouldn't it be 2x\[^{2}\] + 0 =9

OpenStudy (anonymous):

that's what I thought @danirubi

OpenStudy (anonymous):

wait so would it be 2x^2=9 and like im not sure what else I have to do

OpenStudy (anonymous):

well I'm not sure it's 100% correct but I did my best 2(2.25 x 2.25)=9 2(4.5)=9 9=9 x=2.25 hope that makes sense

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

welcome :)